# Proving the existence of localized patterns and saddle node bifurcations in 1D activator-inhibitor type models

### v1

@@ 1 Introduction
where $\delta$, $c$, $d$, $h$, $j$, $a$, and $b$ are positive parameters, and $\Delta$ is the Laplacian operator. These equations arise naturally in various fields such as biology and chemistry, where they model a chemical reaction between two substances $U$ and $V$. The first species is the activator ($U$) with concentration $u$ and the other one is an inhibitor ($V$) with concentration $v$ GrayScott1985 . This system is set in the regime $0<\delta\ll 1$, where the inhibitor diffuses more than the activator. This class of PDEs encompasses many reaction-diffusion systems that have been studied extensively in the literature, such as the Schnakenberg model schankenberg_og ($c=1,d=h=j=0$), the Sel’kov Schnakenberg model selkov_schankenberg_og ($c=1,d=j,h=0$), the Brusselator model brusselator_og (h=c1>0,d=j,h=0h=c-1>0,d=j,h=0), the Root hair model root_hair_og ($a=0,d=j,h>c>0$), and the Glycolysis model glycolysis_og ($c=1,d=j,a=h=0$).
@@ 1.1 Stationary localized solutions
The study of pattern formation originates in the seminal work of Alan Turing turing_morphogenesis . In this paper, Turing demonstrated how patterns can emerge from perturbations of a steady-state solution through what is now referred to as a Turing instability, or diffusion-driven instability. In one spatial dimension, this mechanism typically manifests as a supercritical pitchfork bifurcation from the steady-state solution Meron2015 .
@@
Reaction–Diffusion (RD) systems have attracted significant interest due to their physical relevance, with applications spanning biology root_hair_og ; selkov_schankenberg_og , chemistry GrayScott1985 ; schankenberg_og ; brusselator_og , and ecology SITEUR201481 ; ZELNIK201727 . These systems are particularly compelling because they often exhibit rich and complex behaviors while maintaining a relatively simple mathematical structure. Their nonlinearities are typically polynomial, as in (1), which includes a simple auto-catalytic term of the form $\pm u^{2}v$ in each equation. RD systems have been studied extensively using a variety of analytical approaches. For example, geometric perturbation theory and spatial dynamics have been employed to establish the existence of localized and periodic solutions on long one-dimensional domains in the Gray–Scott model Doelman1997 , later extended to higher dimensions MuratovOsipov2000 and to more general RD systems doelman_geometric . The stability of localized patterns can be investigated using asymptotic expansions, wherein the solution is analyzed through an inner expansion near the pattern and a far-field expansion in the tail. Matching these expansions provides approximations of the eigenvalues of the linearized PDE operator, yielding stability criteria. This method has been successfully applied in the one-dimensional Gierer–Meinhardt model iron_semistrong , the Schnakenberg model Iron2004 , and the Gray–Scott model muratov_semistrong . Another central phenomenon in RD systems is the connection between localized pattern formation and homoclinic snaking arising from a Hamiltonian–Hopf bifurcation, which has been investigated in knobloch_dissipative ; VERSCHUEREN2021132858 .
@@
In recent work, Champneys et al.Champneys_Bistability outlined several mechanisms for pattern formation in RD systems. They investigated the bifurcation structure of the Brusselator model, constructing approximate solutions via matched asymptotic expansions and numerically continuing them to reveal homoclinic snaking. A similar analysis was carried out for the Schnackenberg model in alsaadi_spikes , where the same approach was employed to build approximate solutions and to numerically observe the system’s snaking behavior, alongside a stability analysis of spike solutions. It is important to note that the validity of matched asymptotic expansions is restricted to specific parameter regimes: for the Brusselator, this requires $\delta\ll 1$, while for the Schnackenberg model the appropriate scaling is $a=\epsilon\alpha$, $b=\epsilon^{2}\beta$, with $\epsilon\ll 1$. Building on these results, AlSaadi_unifying_framework extended the analysis to demonstrate the ubiquity of such behavior in systems of the form (1). Using weakly nonlinear analysis, they characterized the bifurcation structure of (1), identifying parameter regions where homoclinic snaking is expected to occur for both large and small parameter values. They then applied matched asymptotic expansions to numerically explore the snaking region and further examined the stability of solutions within this region as the diffusion coefficient $\delta$ increases. It should be emphasized, however, that while their theoretical analysis is posed on $\mathbb{R}$, the corresponding numerical results were obtained on long one-dimensional domains.
@@
In their conclusion, the authors of AlSaadi_unifying_framework raise the critical open question of providing a rigorous proof for their results. Addressing this challenge is essential, as it bridges the gap between formal asymptotics, numerical exploration, and mathematical rigor. In the next subsection, we present a computer-assisted approach that delivers such a rigorous justification, thereby validating and strengthening their numerical observations. Constructive existence proofs of patterns, similar to Figure 1, as well as a constructive proof of a saddle node will be provided. In particular, our method is based on the computer-assisted analysis presented in gs_cadiot_blanco . As these models arise from physical context, the question of the stability of these solutions is also studied by rigorously controlling the spectrum of the jacobian of the linearization around an approximate solution using the approach of cadiot2025stabilityanalysislocalizedsolutions
@@ 1.2 A Computer-assisted approach
One way to approach solving semilinear PDEs on $\mathbb{R}^{m}$ is to use computer-assisted proofs (CAPs) techniques. CAPs can help derive a constructive methodology that relies on a Newton-Kantorovich argument. It involves the computation of specific bounds, which can be evaluated rigorously on the computer by the use of interval arithmetic for instance julia_interval . For the ODE case, one might use the parameterization method MR1976079 ; MR1976080 ; MR2177465 . This method relies on the rigorous parametrization of invariant manifolds at the zero solution. Then, this allows one to formulate a projected boundary value problem which one can treat using Chebyshev series or splines. Still for ODEs, the analysis derived in breden2025solutionsdifferentialequationsfreudweighted ; breden2024constructiveproofssemilinearpdes allows to produce constructive existence proofs in a well-chosen class of equations. Indeed, such equations possess a confining potential, allowing the resolvent of the differential operator to be compact on a well-chosen space of functions. Note that the method derived in breden2024constructiveproofssemilinearpdes also applies to PDEs possessing a linear part of the form $\Delta u+\frac{x}{2}\cdot\nabla u$. Moving to PDEs, but in the weak formulation, the authors of plum_numerical_verif presented a method for proving weak solutions to second and fourth-order PDEs. Their approach relies on the rigorous control of the spectrum of the linearization around an approximate solution. Then, they use a homotopy argument and the Temple-Lehmann-Goerisch method (see Section 10.2 in plum_numerical_verif ). This approach was then applied by Wunderlich in plum_thesis_navierstokes . Wunderlich proved the existence of a weak solution to the Navier-Stokes equations defined on an infinite strip with an obstacle. As far as strong solutions are concerned, olivier_radial provides a methodology for proving the existence of radially symmetric solutions. Using the radially symmetric ansatz, the PDE is transformed into an ordinary differential equation. The approach then relies on a rigorous enclosure of the center stable manifold by using a Lyapunov-Perron operator. This allows one to solve a boundary value problem on $(0,\infty)$. Removing the radially symmetric assumption, the authors of unbounded_domain_cadiot provide a general method for proving the existence and local uniqueness of localized solutions to semilinear autonomous PDEs on $\mathbb{R}^{m}$. The approach is based on Fourier series and describes the necessary tools to construct an approximate solution, $u_{0}$, and an approximate inverse $\mathbb{A}$ of the linearization $D\mathbb{F}(u_{0})$ about $u_{0}$. An application to the 1D Kawahara equation was provided in the same paper. In sh_cadiot , the method of unbounded_domain_cadiot was applied to the planar Swift Hohenberg PDE. The method was then extended further by the authors of gs_cadiot_blanco . In unbounded_domain_cadiot , the method was presented for scalar PDEs. In gs_cadiot_blanco , the authors generalized the approach to systems of PDEs. This allows for rigorous proofs of localized solutions to systems of PDEs. The approach was demonstrated on the 2D Gray-Scott system of equations. Then, in symmetry_blanco_cadiot , the authors extended further this methodology to treat existence proofs of symmetric solutions, where the symmetry might be given by a general group of symmetries. The 2D planar Swift Hohenberg equation was again used as an illustration, allowing for the authors to obtain existence proofs of solutions with various dihedral symmetries.
@@
In this paper, we base our analysis on the method developed in gs_cadiot_blanco . Our goal is to provide rigorous validations to some of the observations made by the authors of AlSaadi_unifying_framework . In fact, we aim at applying similar ideas as in gs_cadiot_blanco to a well-chosen augmented system in order to prove the existence of saddle node bifurcations. This provides a novel computer-assisted approach for proving the existence of saddle-node bifurcations for localized solutions.
@@
In this paper, we will focus our attention on stationary localized solutions, that is $\mathbf{u}:\mathbb{R}\to\mathbb{R}^{2}$ such that $\mathbb{f}(\mathbf{u})=0$ and satisfying $\mathbf{u}(x)\to 0$ as $|x|\to\infty.$ Given a well-chosen Hilbert space $\mathcal{H}_{e}$ (cf. Section 2), containing even localized functions, we look for zeros of $\mathbb{f}$ in $\mathcal{H}_{e}$. In fact, we demonstrate that $\mathbb{f}:\mathcal{H}_{e}\to L^{2}_{e}$ is a smooth operator, where $L^{2}_{e}$ is the restriction of even functions of $L^{2}(\mathbb{R})\times L^{2}(\mathbb{R})$. Then, following the analysis derived in unbounded_domain_cadiot ; gs_cadiot_blanco , we prove the existence of zeros of $\mathbb{f}$ using a Newton-Kantorovich approach (cf. Section 3). In fact, our analysis relies on the construction of an approximate solution $\bar{\mathbf{u}}\in\mathcal{H}_{e}$ and of an approximate inverse $\mathbb{A}:{L}^{2}_{e}\to\mathcal{H}_{e}$ for the Fréchet derivative $D\mathbb{f}(\bar{\mathbf{u}})$. Indeed, we can then define a Newton-like fixed point operator $\mathbb{T}$ given as
@@
and some extra non-degeneracy conditions (given in jp_saddle_node for instance). In particular, jp_saddle_node provides an equivalent definition of a saddle-node bifurcation, thanks to isolated zeros of a well-chosen map. In this paper, we consider such a definition, which we explicit below. We introduce the Hilbert spaces $H_{1}=\mathbb{R}\times\mathcal{H}_{e}\times\mathcal{H}_{e}$ and $H_{2}=\mathbb{R}\times L_{e}^{2}\times L^{2}_{2}$ and we define the so-called Saddle-Node Map $\mathbb{F}:H_{1}\to H_{2}$ as follows
@@
and where $\alpha\in\mathbb{R}$, $\phi:\mathcal{H}_{e}\to\mathbb{R}$ is a non-trivial linear functional. Then, using jp_saddle_node , we have the following definition for a saddle-node bifurcation.
@@ 2 Setup of the Problem
Adopting the notation of unbounded_domain_cadiot , we look for $\mathbf{u}=(u_{1},u_{2})$ such that
@@
for all $\xi\in\mathbb{R}$. We now recall a necessary assumption from gs_cadiot_blanco that our framework requires. We will later verify it for a class of Activator-Inhibitor models (2).
@@
for all $\mathbf{u},\mathbf{v}\in\mathcal{H}.$ In particular, note that $\mathbb{l}:\mathcal{H}\to L^{2}$ is a well-defined bounded linear operator, which is actually an isometric isomorphism. We now want to prove that the operator $\mathbb{g}:\mathcal{H}\to L^{2}$ is a smooth operator. For this matter, we recall two spaces from gs_cadiot_blanco : $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ given as
@@
In fact, using $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, gs_cadiot_blanco provides sufficient conditions for which products on $\mathcal{H}\times\mathcal{H}\to L^{2}$ are well-defined. Using this result, we obtain the following lemma.
@@
where we used Plancherel’s theorem. But now, using the proof of Lemma 2.1 in gs_cadiot_blanco , we get
@@
Finally, let us analyze $\|\hat{\mathbf{u}}\|_{1}$. Using the proof of Lemma 2.1 in gs_cadiot_blanco again, we get
@@
Using Lemma 2.2, we obtain that $\mathbb{g}:\mathcal{H}\to L^{2}$ is a smooth operator. In particular, this yields that $\mathbb{f}:\mathcal{H}\to L^{2}$ is smooth as well. In fact, Proposition 2.1 in gs_cadiot_blanco yields that solutions to (2) in $\mathcal{H}$ are actually infinitely differentiable, meaning that our analysis on $\mathcal{H}$ will yield strong solutions to (2).

### v2

@@ 1 Introduction
where $\delta$, $c$, $d$, $h$, $j$, $a$, and $b$ are positive parameters, and $\Delta$ is the Laplacian operator. These equations arise naturally in various fields such as biology and chemistry, where they model a chemical reaction between two substances $U$ and $V$. The first species is the activator ($U$) with concentration $u$ and the other one is an inhibitor ($V$) with concentration $v$ Gray and Scott (1985). This system is set in the regime $0<\delta\ll 1$, where the inhibitor diffuses more than the activator. This class of PDEs encompasses many reaction-diffusion systems that have been studied extensively in the literature, such as the Schnakenberg model Schnakenberg (1979) ($c=1,d=h=j=0$), the Sel’kov Schnakenberg model Sel’kov (1968) ($c=1,d=j,h=0$), the Brusselator model Prigogine and Lefever (1968)(h=c1>0,d=j,h=0h=c-1>0,d=j,h=0), the Root hair model Payne and Grierson (2009)($a=0,d=j,h>c>0$), and the Glycolysis model Tyson and Kauffman (1975)($c=1,d=j,a=h=0$).
@@ 1.1 Stationary localized solutions
The study of pattern formation originates in the seminal work of Alan Turing Turing (1952). In this paper, Turing demonstrated how patterns can emerge from perturbations of a steady-state solution through what is now referred to as a Turing instability, or diffusion-driven instability. In one spatial dimension, this mechanism typically manifests as a supercritical pitchfork bifurcation from the steady-state solution Meron (2015).
@@
Reaction–Diffusion (RD) systems have attracted significant interest due to their physical relevance, with applications spanning biology Payne and Grierson (2009); Sel’kov (1968), chemistry Gray and Scott (1985); Schnakenberg (1979); Prigogine and Lefever (1968), and ecology Siteur et al. (2014); Zelnik et al. (2017). These systems are particularly compelling because they often exhibit rich and complex behaviors while maintaining a relatively simple mathematical structure. Their nonlinearities are typically polynomial, as in (1), which includes a simple auto-catalytic term of the form $\pm u^{2}v$ in each equation. RD systems have been studied extensively using a variety of analytical approaches. For example, geometric perturbation theory and spatial dynamics have been employed to establish the existence of localized and periodic solutions on long one-dimensional domains in the Gray–Scott model Doelman et al. (1997), later extended to higher dimensions Muratov and Osipov (2000) and to more general RD systems Doelman and Veerman (2015). The stability of localized patterns can be investigated using asymptotic expansions, wherein the solution is analyzed through an inner expansion near the pattern and a far-field expansion in the tail. Matching these expansions provides approximations of the eigenvalues of the linearized PDE operator, yielding stability criteria. This method has been successfully applied in the one-dimensional Gierer–Meinhardt model Iron et al. (2001), the Schnakenberg model Iron et al. (2004), and the Gray–Scott model Muratov and Osipov (2002). Another central phenomenon in RD systems is the connection between localized pattern formation and homoclinic snaking arising from a Hamiltonian–Hopf bifurcation, which has been investigated in Knobloch (2015); Verschueren and Champneys (2021).
@@
In recent work, Champneys et al.Champneys et al. (2021) outlined several mechanisms for pattern formation in RD systems. They investigated the bifurcation structure of the Brusselator model, constructing approximate solutions via matched asymptotic expansions and numerically continuing them to reveal homoclinic snaking. A similar analysis was carried out for the Schnackenberg model in Al saadi et al. (2021), where the same approach was employed to build approximate solutions and to numerically observe the system’s snaking behavior, alongside a stability analysis of spike solutions. It is important to note that the validity of matched asymptotic expansions is restricted to specific parameter regimes: for the Brusselator, this requires $\delta\ll 1$, while for the Schnackenberg model the appropriate scaling is $a=\epsilon\alpha$, $b=\epsilon^{2}\beta$, with $\epsilon\ll 1$. Building on these results, Al Saadi et al. (2021) extended the analysis to demonstrate the ubiquity of such behavior in systems of the form (1). Using weakly nonlinear analysis, they characterized the bifurcation structure of (1), identifying parameter regions where homoclinic snaking is expected to occur for both large and small parameter values. They then applied matched asymptotic expansions to numerically explore the snaking region and further examined the stability of solutions within this region as the diffusion coefficient $\delta$ increases. It should be emphasized, however, that while their theoretical analysis is posed on $\mathbb{R}$, the corresponding numerical results were obtained on long one-dimensional domains.
@@
In their conclusion, the authors of Al Saadi et al. (2021) raise the critical open question of providing a rigorous proof for their results. Addressing this challenge is essential, as it bridges the gap between formal asymptotics, numerical exploration, and mathematical rigor. In the next subsection, we present a computer-assisted approach that delivers such a rigorous justification, thereby validating and strengthening their numerical observations. Constructive existence proofs of patterns, similar to Figure 1, as well as a constructive proof of a saddle node will be provided. In particular, our method is based on the computer-assisted analysis presented in Cadiot and Blanco (2025). As these models arise from physical context, the question of the stability of these solutions is also studied by rigorously controlling the spectrum of the jacobian of the linearization around an approximate solution using the approach of Cadiot (2025b)
@@ 1.2 A Computer-assisted approach
One way to approach solving semilinear PDEs on $\mathbb{R}^{m}$ is to use computer-assisted proofs (CAPs) techniques. CAPs can help derive a constructive methodology that relies on a Newton-Kantorovich argument. It involves the computation of specific bounds, which can be evaluated rigorously on the computer by the use of interval arithmetic for instance Benet and Sanders (2022). For the ODE case, one might use the parameterization method Cabré et al. (2003a, b, 2005). This method relies on the rigorous parametrization of invariant manifolds at the zero solution. Then, this allows one to formulate a projected boundary value problem which one can treat using Chebyshev series or splines. Still for ODEs, the analysis derived in Breden and Chu (2025, 2024) allows to produce constructive existence proofs in a well-chosen class of equations. Indeed, such equations possess a confining potential, allowing the resolvent of the differential operator to be compact on a well-chosen space of functions. Note that the method derived in Breden and Chu (2024) also applies to PDEs possessing a linear part of the form $\Delta u+\frac{x}{2}\cdot\nabla u$. Moving to PDEs, but in the weak formulation, the authors of Nakao et al. (2019) presented a method for proving weak solutions to second and fourth-order PDEs. Their approach relies on the rigorous control of the spectrum of the linearization around an approximate solution. Then, they use a homotopy argument and the Temple-Lehmann-Goerisch method (see Section 10.2 in Nakao et al. (2019)). This approach was then applied by Wunderlich in Wunderlich (2022). Wunderlich proved the existence of a weak solution to the Navier-Stokes equations defined on an infinite strip with an obstacle. As far as strong solutions are concerned, van den Berg et al. (2023) provides a methodology for proving the existence of radially symmetric solutions. Using the radially symmetric ansatz, the PDE is transformed into an ordinary differential equation. The approach then relies on a rigorous enclosure of the center stable manifold by using a Lyapunov-Perron operator. This allows one to solve a boundary value problem on $(0,\infty)$. Removing the radially symmetric assumption, the authors of Cadiot et al. (2024) provide a general method for proving the existence and local uniqueness of localized solutions to semilinear autonomous PDEs on $\mathbb{R}^{m}$. The approach is based on Fourier series and describes the necessary tools to construct an approximate solution, $u_{0}$, and an approximate inverse $\mathbb{A}$ of the linearization $D\mathbb{F}(u_{0})$ about $u_{0}$. An application to the 1D Kawahara equation was provided in the same paper. In Cadiot et al. (2025), the method of Cadiot et al. (2024) was applied to the planar Swift Hohenberg PDE. The method was then extended further by the authors of Cadiot and Blanco (2025). In Cadiot et al. (2024), the method was presented for scalar PDEs. In Cadiot and Blanco (2025), the authors generalized the approach to systems of PDEs. This allows for rigorous proofs of localized solutions to systems of PDEs. The approach was demonstrated on the 2D Gray-Scott system of equations. Then, in Blanco and Cadiot (2025b), the authors extended further this methodology to treat existence proofs of symmetric solutions, where the symmetry might be given by a general group of symmetries. The 2D planar Swift Hohenberg equation was again used as an illustration, allowing for the authors to obtain existence proofs of solutions with various dihedral symmetries.
@@
In this paper, we base our analysis on the method developed in Cadiot and Blanco (2025). Our goal is to provide rigorous validations to some of the observations made by the authors of Al Saadi et al. (2021). In fact, we aim at applying similar ideas as in Cadiot and Blanco (2025) to a well-chosen augmented system in order to prove the existence of saddle node bifurcations. This provides a novel computer-assisted approach for proving the existence of saddle-node bifurcations for localized solutions.
@@
In this paper, we will focus our attention on stationary localized solutions, that is $\mathbf{u}:\mathbb{R}\to\mathbb{R}^{2}$ such that $\mathbb{f}(\mathbf{u})=0$ and satisfying $\mathbf{u}(x)\to 0$ as $|x|\to\infty.$ Given a well-chosen Hilbert space $\mathcal{H}_{e}$ (cf. Section 2), containing even localized functions, we look for zeros of $\mathbb{f}$ in $\mathcal{H}_{e}$. In fact, we demonstrate that $\mathbb{f}:\mathcal{H}_{e}\to L^{2}_{e}$ is a smooth operator, where $L^{2}_{e}$ is the restriction of even functions of $L^{2}(\mathbb{R})\times L^{2}(\mathbb{R})$. Then, following the analysis derived in Cadiot et al. (2024); Cadiot and Blanco (2025), we prove the existence of zeros of $\mathbb{f}$ using a Newton-Kantorovich approach (cf. Section 3). In fact, our analysis relies on the construction of an approximate solution $\bar{\mathbf{u}}\in\mathcal{H}_{e}$ and of an approximate inverse $\mathbb{A}:{L}^{2}_{e}\to\mathcal{H}_{e}$ for the Fréchet derivative $D\mathbb{f}(\bar{\mathbf{u}})$. Indeed, we can then define a Newton-like fixed point operator $\mathbb{T}$ given as
@@
and some extra non-degeneracy conditions (given in Lessard et al. (2017) for instance). In particular, Lessard et al. (2017) provides an equivalent definition of a saddle-node bifurcation, thanks to isolated zeros of a well-chosen map. In this paper, we consider such a definition, which we explicit below. We introduce the Hilbert spaces $H_{1}=\mathbb{R}\times\mathcal{H}_{e}\times\mathcal{H}_{e}$ and $H_{2}=\mathbb{R}\times L_{e}^{2}\times L^{2}_{2}$ and we define the so-called Saddle-Node Map $\mathbb{F}:H_{1}\to H_{2}$ as follows
@@
and where $\alpha\in\mathbb{R}$, $\phi:\mathcal{H}_{e}\to\mathbb{R}$ is a non-trivial linear functional. Then, using Lessard et al. (2017), we have the following definition for a saddle-node bifurcation.
@@ 2 Setup of the Problem
Adopting the notation of Cadiot et al. (2024), we look for $\mathbf{u}=(u_{1},u_{2})$ such that
@@
for all $\xi\in\mathbb{R}$. We now recall a necessary assumption from Cadiot and Blanco (2025) that our framework requires. We will later verify it for a class of Activator-Inhibitor models (2).
@@
for all $\mathbf{u},\mathbf{v}\in\mathcal{H}.$ In particular, note that $\mathbb{l}:\mathcal{H}\to L^{2}$ is a well-defined bounded linear operator, which is actually an isometric isomorphism. We now want to prove that the operator $\mathbb{g}:\mathcal{H}\to L^{2}$ is a smooth operator. For this matter, we recall two spaces from Cadiot and Blanco (2025) : $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ given as
@@
In fact, using $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, Cadiot and Blanco (2025) provides sufficient conditions for which products on $\mathcal{H}\times\mathcal{H}\to L^{2}$ are well-defined. Using this result, we obtain the following lemma.
@@
where we used Plancherel’s theorem. But now, using the proof of Lemma 2.1 in Cadiot and Blanco (2025), we get
@@
Finally, let us analyze $\|\hat{\mathbf{u}}\|_{1}$. Using the proof of Lemma 2.1 in Cadiot and Blanco (2025) again, we get
@@
Using Lemma 2.2, we obtain that $\mathbb{g}:\mathcal{H}\to L^{2}$ is a smooth operator. In particular, this yields that $\mathbb{f}:\mathcal{H}\to L^{2}$ is smooth as well. In fact, Proposition 2.1 in Cadiot and Blanco (2025) yields that solutions to (2) in $\mathcal{H}$ are actually infinitely differentiable, meaning that our analysis on $\mathcal{H}$ will yield strong solutions to (2).